the chemist used a solution of iron nitrate, Fe(NO3)3 with a concentration of 0.010 mol l–1. If they were preparing one litre of this solution, what mass of iron nitrate should they weigh out?

For your calculations use the following relative atomic mass (RAM) values:

RAM (Fe) = 56

RAM (N) = 14

RAM (O) = 16


1
Expert's answer
2021-01-12T07:03:36-0500

compound = "Fe(NO_3)_2"

molar mass of compound = Fe + 2N + 6O = 56 + 2(14) + 6(16) = 180g/mol


concentration = 0.010 mol L-1


If the volume of the solution is 1L, the amount of Iron Nitrate = c × V = 0.010 × 1 = 0.010 mol



Since 1 mole of Iron Nitrate = 180g

0.010 mol = x


x = "\dfrac{0.010×180}{1}" = 1.80g


Therefore, If they were preparing one litre of the solution, 1.80g of iron nitrate should be weighed out.